OpenStudy (anonymous):

If $2000 is invested at an interest rate of 3.5% per year, compounded continuously, find the value of the investment after the given number of years: a) 2 years b) 4 years c) 12 years

OpenStudy (anonymous):

\[2000\times e^{2\times .035}\] \[2000\times e^{4\times .035}\] \[2000\times e^{12\times .035}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw